OpenStudy (anonymous):

Express y in terms of x ln y = 4 ln x + ln 9 how do i do that?

OpenStudy (anonymous):

This seems incorrectly typed.

OpenStudy (anonymous):

it is not

OpenStudy (mathstudent55):

\(\ln y = 4 \ln x + \ln 9\) \(\ln y = \ln x^4 + \ln 9\) \(\ln y = \ln 9x^4 \) \(y = 9x^4 \)